# Geometric Transformations on Null Curves in the Anti-de Sitter $3$-Space

### v1

@@ 5. Construction Procedure
Refer to captionRefer to captionRefer to caption
Figure 9. Left: The null curve $s\in\mathbb{R}\longmapsto\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\left(\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma(s,0))\right)$. Center: The two disjoint, but congruent torus knots. Right: The arc connecting the two torus knots. For this example, we have fixed $m=4$, $n=1$, $p=1.4$ and $r=1$.
Notice
The present paper is a continuation of the work [8], which is submitted for a possible publication but it still in the review process. As a consequence, this work will not be submitted to a journal until [8] is already accepted/published.

### v2

@@ 5 Construction procedure
Figure 9: Left: The null curve $s\in\mathbb{R}\longmapsto\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\bigl{(}\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma(s,0))\bigr{)}$. Center: The two disjoint, but congruent torus knots. Right: The arc connecting the two torus knots. For this example, we have fixed $m=4$, $n=1$, $p=1.4$ and $r=1$.
Acknowledgments
Authors partially supported by PRIN 2017 and PRIN 2022 “Real and Complex Manifolds: Topology, Geometry and Holomorphic Dynamics” (prott. 2017JZ2SW5-004 and 2022AP8HZ9-003); and by the GNSAGA of INdAM. The authors gratefully acknowledge the warm hospitality of the Department of Mathematics at the Politecnico di Torino and of the Texas Tech University Center in Sevilla. The authors would like to thank the referees for carefully reviewing the paper.